What is the weight percent (w/w) of NaOH in a solution prepared by dissolving 30.0g of NaOH in 120.0g of H2O?




1
Expert's answer
2022-05-25T02:47:01-0400

weight percent -  grams of solute/(grams of solute plus solvent)*100

w/w=(30/(30+120))*100=20%